### v1

# Analytic Langlands correspondence from SoV

Abstract
The analytic Langlands correspondence proposed by Etingof, Frenkel and Kazhdan describes the solution to the spectral problems naturally arising in the quantisation of the Hitchin integrable systems in terms of real opers, certain second order differential operators on a Riemann surface having real monodromy. We prove this correspondence in the cases associated to the group $\mathrm{PSL}(2,\mathbb{C})$, and Riemann surfaces of genus zero with a number of punctures larger than three. A crucial ingredient is a unitary integral transformation mapping products of solutions to the ordinary differential equation associated to a real oper to eigenfunctions of the quantised Hitchin Hamiltonians. This allows us to construct joint eigenfunctions of Hecke operators and Hitchin Hamiltonians from real opers.
1 Introduction
The analytic Langlands correspondence has been introduced in [EFK1]. It supplements the geometric Langlands correspondence by analytic aspects, claiming a correspondence between square-integrable eigenfunctions of the Hamiltonians of the quantised Hitchin integrable system on a Riemann surface $C$, and real opers, a certain class of flat connections on $C$ having real holonomy. It refines the correspondence which had previously been proposed in [Te] between single-valued eigenfunctions of the Hamiltonians of the Hitchin integrable system, and real opers. Such correspondences have interesting relations outlined in [GW, GT] to supersymmetric quantum field theories, and to two-dimensional conformal field theories.
The proposal made in [EFK1] has two main ingredients helping to gain analytic control. The first is to work in the Hilbert space framework offered by the natural scalar product on spaces of half-densities on the moduli spaces $\mathrm{Bun}_{G}(C)$ of holomorphic $G$-bundles on $C$. Another crucial ingredient is the use of the Hecke-operators, natural families of integral operators on the spaces of half-densities on $\mathrm{Bun}_{G}(C)$ [EFK2]. As eigenfunctions of the Hecke-operators turn out to be simultaneously eigenfunctions of the Hitchin Hamiltonians, one may start by studying the spectral problem for the Hecke operators. This spectral problem is more easily seen to be well-defined, as the Hecke operators are compact self-adjoint operators, while the Hamiltonians of the quantised Hitchin system are unbounded.
A proof of the analytic Langlands correspondence so far only appears to be available in a few cases, associated to $G=\mathrm{SL}(2)$, and Riemann surfaces $C=C_{g,n}$ of genus $g=0$ and $n=4,5$ punctures [EFK1, EFK3]. The goal of this paper is to offer a proof for arbitrary $n$.
To this aim we shall adopt a strategy known as Separation of Variables ($\mathsf{SoV}$) in the literature on quantum integrable models. It amounts to the construction of a unitary integral transformation mapping the product of solutions to an ordinary differential equation to a solution of the eigenvalue equations of our interest. The ordinary differential equations appearing in this context are closely related to the real opers appearing on one side of the analytic Langlands correspondence. The $\mathsf{SoV}$ method is powerful as it allows one to construct eigenfunctions of the Hitchin Hamiltonians from real opers. One may thereby prove that the condition to correspond to a real oper is not only necessary, but also sufficient for being an eigenfunction.
To carry this program out in practise, it will be crucial to understand how the Hecke operators interact with the integral transformations defined by the $\mathsf{SoV}$ method. This will allow us to prove that the image of a real oper under the $\mathsf{SoV}$-transformation is an eigenfunction of the Hecke operators. It turns out that the Hecke eigenvalue property follows from the observation that the result of a Hecke-modification at a point $z\in C\equiv C_{0,n}$ can be represented by a $\mathsf{SoV}$-transformation associated to the Riemann surface $C^{\prime}:=C\setminus\{z\}\simeq C_{0,n+1}$. Proving this fact turns out to be the main technical issue arising in this context.
The application of the $\mathsf{SoV}$ method circumvents a difficulty that appears in other approaches to the analytic Langlands correspondence. It originates from the fact that the Hitchin Hamiltonians are singular at the locus in $\mathrm{Bun}_{G}(C)$ associated to the “wobbly” bundles, bundles admitting nilpotent Higgs fields. Checking that the resulting singular behaviour of the eigenfunctions at the locus does not spoil square-integrability can be difficult, having to rely on detailed information on the geometry of the wobbly locus. The $\mathsf{SoV}$ transformation, being unitary, maps the check of square-integrability to the simpler problem to check square-integrability of products of solutions to the oper equation with respect to a suitable measure.
We expect that the strategy used in this paper admits a generalisation to Riemann surfaces $C=C_{g,n}$ of higher genus $g>0$ using the classical version of the $\mathsf{SoV}$-transformation developed in [DT] as geometric groundwork.
2 Analytic Langlands correspondence
We will study the quantised Hitchin systems on $C=\mathbb{P}^{1}\setminus\{z_{1},\dots,z_{n}\}$. In order to formulate the main results it will be convenient to assume, without essential loss of generality, that $z_{n}=\infty$. The relevant geometrical background is described in [EFK3].
2.1 Quantisation of the Hitchin system
We shall define two Hilbert spaces ${\mathcal{H}}_{i}$, $i=0,1$, of homogeneous and translation invariant functions. The elements of ${\mathcal{H}}_{i}$ are represented by functions $\Psi_{i}$ of n1n-1 complex variables $\bm{x}=(x_{1},\dots,x_{n-1})$ satisfying
$\begin{aligned} &\Psi_{i}(x_{1}+\delta,\dots,x_{n-1}+\delta)=\Psi_{i}(x_{1},\dots,x_{n-1}),\\ &\Psi_{i}(\rho x_{1},\dots,\rho x_{n-1})=|\rho|^{2J_{i}}\Psi_{i}(x_{1},\dots,x_{n-1}),\end{aligned},\qquad i=0,1,$ (2.1)
where
$J_{i}=\sum_{r=1}^{n-1}j_{r}-j_{n,i},\qquad j_{n,i}=\bigg\{\begin{aligned} j_{n}\quad\;\;&\text{for}\;\;i=0,\\ -j_{n}-1\;\;&\text{for}\;\;i=1.\end{aligned}$ (2.2)
It is well-known that such functions can be used to represent invariants in tensor products of $n$ spherical principal series representations of $\mathrm{SL}(2,{\mathbb{C}})$ if $j_{r}\in-\frac{1}{2}+\mathrm{i}{\mathbb{R}}$, $r=1,\dots,n$. Translation invariance allows one to represent the functions $\Psi_{i}$ by the homogenous functions $\Psi_{i}(0,x_{2},\dots,x_{n-1})$. Away from $x_{n-1}=0$ one may represent such functions in the form
$\Psi_{i}(0,x_{2},\dots,x_{n-1})=|x_{n-1}|^{2J_{i}}\psi_{i}(x_{2}/x_{n-1},\dots,x_{n-2}/x_{n-1}).$ (2.3)
Similar representations exist away from $x_{r}=0$, $r=2,\dots,n-1$. It follows that the functions $\psi_{i}$ on the right side of (2.3) represent sections of line bundles ${\mathcal{L}}_{J_{i}}$, $i=0,1$, of densities on the complex projective space ${\mathbb{P}}^{d}$, d=n3d=n-3, in a particular chart.
We may observe that the Hilbert space structure on $L^{2}({\mathbb{C}}^{d})$ allows us to define a natural norm on spaces of regular homogenous functions $\Psi$. With the help of (2.3) we may represent this norm as
$\lVert\Psi\rVert^{2}=\int_{{\mathbb{C}}^{d}}d^{2d}\bm{z}\;|\psi(\bm{z})|^{2},\qquad\bm{z}=(z_{2},\dots,z_{n-2}).$ (2.4)
Representatives for the scaling equivalence class can be represented in many other ways than (2.3). When $J\in-\frac{n-2}{2}+\mathrm{i}{\mathbb{R}}$ it can be checked that $\lVert\Psi\rVert^{2}$ does not depend on the representation, explaining why (2.4) defines a natural Hilbert space ${\mathcal{H}}_{J}$ of square-integrable sections of ${\mathcal{L}}_{J}$.
We furthermore note that the Hilbert spaces ${\mathcal{H}}_{i}$, $i=0,1$, are canonically isomorphic, with unitary isomorphisms $\mathsf{H}_{\infty}:{\mathcal{H}}_{i}\to{\mathcal{H}}_{i+1}$ being defined by sending $\psi_{i}$ to $\psi_{i+1}$, for $i\in{\mathbb{Z}}/2$, but the representation in terms of homogenous functions differs.
The quantum integrable structure of the Hitchin system is defined by the Hamiltonians. In order to define the Hamiltonians, we shall first introduce the differential operators
${\mathcal{D}}_{r}^{+}=\partial_{x_{r}},\qquad{\mathcal{D}}_{r}^{0}=x_{r}\partial_{x_{r}}-j_{r},\qquad{\mathcal{D}}_{r}^{-}=2j_{r}x_{r}-x_{r}^{2}\partial_{x_{r}},$ (2.5)
allowing us to define
$\mathsf{H}_{r}=\sum_{s\neq r}\frac{\eta_{ab}{\mathcal{D}}_{r}^{a}{\mathcal{D}}_{r}^{b}}{z_{r}-z_{s}},\qquad r=1,\dots,n-1,$
where $\eta^{00}=1$, $\eta^{+-}=\eta^{-+}=-\frac{1}{2}$, and $\eta^{ab}=0$ otherwise. One should keep in mind that only n3n-3 out of the n1n-1 operators $\mathsf{H}_{r}$ are linearly independent. We are here considering a slightly more general family of operators compared with [EFK3], having parameters $j_{r}\in-\frac{1}{2}+\mathrm{i}{\mathbb{R}}$ rather than $j_{r}=-\frac{1}{2}$ as considered in [EFK3].
The differential operators ${\mathsf{H}}_{r}$ will be unbounded, as usual. In order to control the functional-analytic aspects it will be advantageous to introduce the so-called Baxter operators.
2.2 Baxter operators
We shall define the Baxter operators $\mathsf{Q}_{y}$ as the composition $\mathsf{Q}_{y}=\mathbf{H}_{\infty}\cdot\mathbf{H}_{y}$, with the Hecke-operators $\mathbf{H}_{y}:{\mathcal{H}}_{i}\to{\mathcal{H}}_{i+1}$, being defined as
$(\mathbb{H}_{y,x}\Psi)(x_{1},\dots,x_{n-1})=\prod_{r=1}^{n-1}\frac{|x_{r}-x|^{4j_{r}}}{|z_{r}-z|^{2j_{r}}}\Psi\big(y_{1}(x_{1},x,y),\dots,y_{n-1}(x_{n-1},x,y)\big),$ (2.6)
$\mathbf{H}_{y}\Psi=\int_{{\mathbb{C}}}d^{2}x\;\mathbb{H}_{y,x}\Psi,\qquad\qquad y_{r}(x_{r},x,y)=-\frac{z_{r}-y}{x_{r}-x}.$
The main properties of operators $\mathsf{Q}_{y}$ have been proven in [EFK3] for $j_{r}=-\frac{1}{2}$, $r=1,\dots,n$. We note that the proofs in [EFK3] can be generalised almost verbatim to the cases with $j_{r}\in-\frac{1}{2}+\mathrm{i}{\mathbb{R}}$ studied here. The Baxter operators are compact, hence bounded, for all $y\in{\mathbb{C}}\setminus\{z_{1},\dots,z_{n-1}\}$. They are furthermore self-adjoint, and mutually commuting,
$\big[\,\mathsf{Q}_{y}\,,\mathsf{Q}_{y^{\prime}}\,\big]=0,\qquad\forall\;y,y^{\prime}\in{\mathbb{C}}\setminus\{z_{1},\dots,z_{n-1}\}.$
The spectral problem for the family of Baxter operators is therefore well-posed, and the abstract spectral theorem ([EFK3, Theorem 3.17]) asserts existence of a discrete set $\mathcal{S}$} such that
${\mathcal{H}}_{i}\simeq\bigoplus_{\chi\in{\mathcal{S}}}{\mathcal{H}}_{\chi},$ (2.7)
where ${\mathcal{H}}_{\chi}$ are the eigenspaces for fixed eigenvalues $\chi$, represented by functions $\chi(y,\bar{y})$ on $C$ such that $\mathsf{Q}_{y}\Psi_{\chi}=\chi(y,\bar{y})\Psi_{\chi}.$ The spaces ${\mathcal{H}}_{\chi}$ are finite-dimensional for all $\chi\in{\mathcal{S}}$.
[EFK3, Proposition 4.3] clarifies the relation between Baxter operators and Hamiltonians, which can be concisely formulated as the operator differential equation
$(\partial_{y}^{2}+\mathsf{T}(y))\mathsf{Q}_{w}=0,\qquad\mathsf{T}(y)=\sum_{r=1}^{n-1}\bigg(\frac{\delta_{r}}{(y-z_{r})^{2}}+\frac{{\mathsf{H}}_{r}}{y-z_{r}}\bigg),$ (2.8)
where $\delta_{r}=j_{r}(j_{r}+1)$. We refer to [EFK3] for the details. This implies that eigenstates of the Baxter operators are eigenstates of ${\mathsf{H}}_{r}$, $r=1,\dots,n-1$.
2.3 Formulation of the analytic Langlands correspondence
The analytic Langlands correspondence describes the spectrum of Hitchin systems in geometric terms. It claims that there is a one-to-one correspondence between eigenstates and real opers.
Definition 1.
An oper on $C_{0,n}=\mathbb{P}^{1}\setminus\{z_{1},\dots,z_{n}\}$ is a differential operator having the form $\partial_{y}^{2}+t(y)$ on ${\mathbb{C}}\setminus\{z_{1},\dots,z_{n}\}$, with $t(y)$ of the form
$t(y)=\sum_{r=1}^{n}\bigg(\frac{\delta_{r}}{(y-z_{r})^{2}}+\frac{E_{r}}{y-z_{r}}\bigg),$ (2.9)
having a finite limit $\lim_{y\to\infty}y^{4}t(y)$. An oper on $C_{0,n}$ is a real oper if it has real monodromy.
An important role will be played by the solutions to the system of differential equations
$(\partial_{y}^{2}+t(y))\chi(y,\bar{y})=0,\qquad(\partial_{\bar{y}}^{2}+\bar{t}(\bar{y}))\chi(y,\bar{y})=0.$ (2.10)
An oper admits solutions $\chi$ that are single-valued on $C_{0,n}$ if and only if it is real. The solution $\chi$ is uniquely determined by the oper $\partial_{y}^{2}+t(y)$, and vice-versa.
The analytic Langlands correspondence is the content of the following theorem.
Theorem 1.
The eigenspaces ${\mathcal{H}}_{\chi}$ are one-dimensional, and the set $\mathcal{S}$} is in bijection to the set $\mathrm{Op}_{{\mathbb{R}}}(C)$ of all real opers on $C$.
3 Construction of the SoV transformation
The main tool for the proof of Theorem 1 will be an integral transformation referred to as Separation of Variables ($\mathsf{SoV}$) transformation that will be defined in this section. It serves to simplify the eigenvalue problems for Hitchin’s Hamiltonians.
Let us define ${\mathcal{H}}_{\rm Skl}=L^{2}({\mathbb{C}}^{n-3},d\mu_{\rm Skl})$, with
$d\mu_{\rm Skl}(\bm{y})=d^{2(n-3)}\bm{y}\,\prod_{\begin{subarray}{c}r,s=1\\ r<s\end{subarray}}^{n-1}|z_{r}-z_{s}|^{2}\prod_{\begin{subarray}{c}k,l=1\\ k<l\end{subarray}}^{n-3}|y_{k}-y_{l}|^{2}\prod_{r=1}^{n-1}\frac{1}{|\kappa_{r}(\bm{y})|^{2}},$ (3.11)
where
$\quad\kappa_{r}(\bm{y}):=\frac{\prod_{k=1}^{n-3}(z_{r}-y_{k})}{\prod_{s\neq r}(z_{r}-z_{s})},\qquad r=1,\dots,n-1.$ (3.12)
[…]
where $\mathsf{F}_{\bm{j},i}$ is a twisted version of the Fourier transformation of homogeneous translation-invariant functions defined below, and $\mathsf{SoV}$ is defined by a change of variables defined below.
3.1 Fourier-transformation
We note that homogenous locally integrable functions $\Psi$ of $d+1$ variables $x_{2},\dots,x_{d+2}$ naturally define homogenous regular distributions $D_{\Psi}$ on the Schwartz space ${\mathcal{S}}({\mathbb{C}}^{d+1})$ by means of
$\langle D_{\Psi},f\rangle=\int_{{\mathbb{C}}^{d+1}}d^{2(d+1)}\bm{x}\;\Psi(\bm{x})f(\bm{x}),\qquad f\in{\mathcal{S}}({\mathbb{C}}^{d+1}).$ (3.13)
The Fourier-transformation of $D_{\Psi}$ can be defined in the distributional sense, yielding a homogenous distribution $\widetilde{D}_{\Psi}$. It turns out that the homogenous distribution $\widetilde{D}_{\Psi}$ is regular, meaning that there exists a a homogenous function $\widetilde{\Psi}$ such that $\widetilde{D}_{\Psi}=D_{\widetilde{\Psi}}$. Let $\eta_{\rho}:{\mathbb{R}}_{+}\to{\mathbb{R}}$ be a family of smooth functions satisfying $\eta_{\rho}(x)=1$ for $x<\rho$ and $\eta_{\rho}(x)=0$ for $x>2\rho$.
Lemma 1.
$\widetilde{\Psi}$ can be represented for smooth homogenous functions $\Psi$ as the limit
$\widetilde{\Psi}({\bm{k}})=\lim_{\rho\to\infty}\int_{{\mathbb{C}}^{d+1}}d^{2(d+1)}{\bm{x}}\;{\Psi}(\bm{x})\,\eta_{\rho}(\lVert{\bm{x}}\rVert)\,\prod_{r=2}^{d+2}{e^{-2\mathrm{i}\,\mathrm{Im}(k_{r}x_{r})}},$ (3.14)
where ${\bm{k}}=(k_{2},\dots,k_{d+2})$. In this case, (3.14) defines a smooth function $\widetilde{\Psi}({\bm{k}})$.
The proof of [GrS, Lemma 1] can easily be adapted to our case where d=n3d=n-3. We note that
$\widetilde{\Psi}(\lambda{\bm{k}})=|\lambda|^{2\bar{J}}\,\widetilde{\Psi}({\bm{k}}),\qquad\bar{J}:=-J-d-1=-\frac{d+1}{2}-iS.$ (3.15)
The function $\widetilde{\Psi}({\bm{k}})$ therefore defines a homogenous regular distribution on ${\mathbb{C}}^{d+1}$, represented by functions $\widetilde{\Psi}(k_{2},\dots,k_{n-1})$ of the form
$\widetilde{\Psi}(k_{2},\dots,k_{n-1})=|k_{n-1}|^{2\bar{J}}\,\widetilde{\psi}(k_{2}/k_{n-1},\dots,k_{n-2}/k_{n-1}).$ (3.16)
It follows from Lemma (1) that the Fourier-transformation maps the space ${\mathcal{S}}_{J}$ of smooth sections of ${\mathcal{L}}_{J}$ to the space ${\mathcal{S}}_{-J-d-1}$. The definition of the norm given in (2.4) can be applied to $\widetilde{\Psi}$.
Theorem 2.
The Fourier transformation sending $\psi\in{\mathcal{S}}_{J}$ to $\widetilde{\psi}\in{\mathcal{S}}_{-J-d-1}$ is unitary.
The proof in [GrS, Section 3] can be adapted to our case. This implies that the maps $\mathsf{F}_{\bm{j},i}$ have natural extensions from ${\mathcal{S}}_{J_{i}}$ to the spaces ${\mathcal{H}}_{J_{i}}$ of square-integrable sections of ${\mathcal{L}}_{J_{i}}$.
3.2 Separation of Variables
It will be convenient to use “twisted” versions of the Fourier-transformation, sending $\psi$ to
$(\mathsf{F}_{\bm{j},i}\Psi)(\bm{k}):=|\rho(\bm{k})|^{-2(j_{n,i}+1)}\prod_{r=1}^{n-1}|k_{r}|^{2j_{r}+2}\,\widetilde{\Psi}(\bm{k}),\qquad\rho(\bm{k}):=\sum_{r=1}^{n-1}z_{r}k_{r},$ (3.17)
for $i=0,1$, using the notations $j_{n,i}$ introduced in (2.2). The change of variables defined by
$k_{r}(\bm{y},\rho):=\rho\kappa_{r}(\bm{y}),\qquad\kappa_{r}(\bm{y}):=\frac{\prod_{k=1}^{n-3}(z_{r}-y_{k})}{\prod_{s\neq r}(z_{r}-z_{s})},\qquad r=1,\dots,n-1,$ (3.18)
allows us to define
$\Phi(\bm{y},\rho):=(\mathsf{F}_{\bm{j},i}{\Psi})\big(\bm{k}(\bm{y},\rho)\big),$ (3.19)
It follows easily from (3.15) that $\Phi(\bm{y},\lambda\rho)=\Phi(\bm{y},\rho)$, implying that $\Phi(\bm{y},\rho)$ does not depend on $\rho$, defining a function $\phi(\bm{y})$. We then have
Lemma 2.
The map $\widetilde{{\mathcal{H}}}_{-J-d-1}\to{\mathcal{H}}_{\rm Skl}$ sending $\widetilde{\psi}$ to the function $\phi$ defined above is unitary.
Proof.
We note that the change of variables defined by
$\lambda_{r}(\bm{y}):=\frac{\kappa_{r}(\bm{y})}{\kappa_{n-1}(\bm{y})},\quad r=2,\dots,n-2,$ (3.20)
induces the following change of measures
$d^{2(n-3)}\bm{\lambda}=\frac{d^{2(n-3)}\bm{y}}{|\kappa_{n-1}(\bm{y})|^{2(n-2)}}\prod_{\begin{subarray}{c}r,s=1\\ r<s\end{subarray}}^{n-1}|z_{r}-z_{s}|^{2}\prod_{\begin{subarray}{c}k,l=1\\ k<l\end{subarray}}^{n-3}|y_{k}-y_{l}|^{2}.$ (3.21)
It furthermore follows from (3.19) that
$\big|\psi\big(\bm{\lambda}(\bm{y})\big)\big|^{2}=|\phi(\bm{y})|^{2}\,{|\kappa_{n-1}(\bm{y})|^{2(n-2)}}\prod_{r=1}^{n-1}\frac{1}{|\kappa_{r}(\bm{y})|^{2}}.$ (3.22)
By combining (3.21) and (3.22) it follows that
$\int_{{\mathbb{C}}^{n-3}}d^{2(n-3)}\bm{\lambda}\;|\psi(\bm{\lambda})|^{2}=\int_{{\mathbb{C}}^{n-3}}d\mu_{\rm Skl}(\bm{y})\;|\phi(\bm{y})|^{2},$ (3.23)
which is what we wanted to prove. ∎
Unitarity of the map $\mathsf{SoV}$ follows by combining Theorem 2 and Lemma 2.
For any given differential operator $\mathsf{D}$ of the form
$\mathsf{D}(z)=\sum_{\alpha\in{\mathcal{I}}}f_{\alpha}(z)\mathsf{D}^{(\alpha)},$
with $\{\mathsf{D}^{(\alpha)};\alpha\in{\mathcal{I}}\}$ being a finite set of differential operators in the variables $k_{r}$, and $f_{\alpha}$, $\alpha\in{\mathcal{I}}$ being holomorphic functions on $C$, we may define a differential operator denoted as $\mathsf{D}(y_{k})$ by substituting $z$ by $y_{k}$ “from the left”, and pullback under (3.18), more precisely defined by
${\mathsf{D}(y_{k})}:=\sum_{\alpha\in{\mathcal{I}}}f_{\alpha}(y_{k})\mathsf{D}^{(\alpha)}_{\bm{y},\rho}$, (3.24)
with $\mathsf{D}^{(\alpha)}_{\bm{y},\rho}$, $\alpha\in{\mathcal{I}}$, being the differential operators in the variables $y_{1},\dots,y_{d}$ and $\rho$ associated to $\mathsf{D}^{(\alpha)}$ by the change of variables (3.18).
A crucial property of the map $\mathsf{SoV}$ can now be formulated as follows.
Proposition 1.
The Fourier-transformation $\widetilde{\mathsf{T}}(z):=\mathsf{F}_{\bm{j},i}\cdot\mathsf{T}(z)\cdot\mathsf{F}^{-1}_{\bm{j},i}$ of $\mathsf{T}(z)$ satisfies
$\widetilde{\mathsf{T}}(y_{k})=-\partial^{2}_{y_{k}}.$ (3.25)
By composing (3.25) with the Fourier-transformations $\mathsf{F}_{\bm{j},i}$, it follows from (3.25) that $\mathsf{S}_{\bm{j},i}$ intertwines the eigenvalue equations for the Hamiltonians ${\mathsf{H}}_{r}$ with the system of decoupled equations
$(\partial_{y_{k}}^{2}+t(y_{k}))\Phi(\bm{y})=0,\qquad k=1,\dots,d.$ (3.26)
This fact has been discovered in a closely related context in [Skl]. It motivates the terminology “Separation of Variables”. A self-contained proof is given below.
Proof.
As a preparation, let us note that $\widetilde{\mathsf{T}}(z)$ can be more explicitly represented as
$\widetilde{\mathsf{T}}(z)=-(\mathsf{a}(z))^{2}+\mathsf{a}^{\prime}(z)-\mathsf{c}(z)\mathsf{b}(z),$ (3.27)
where
$\mathsf{c}(z)=\sum_{r=1}^{n-1}\frac{k_{r}}{z-z_{r}},\quad\mathsf{a}(z)=\sum_{r=1}^{n-1}\frac{-k_{r}}{z-z_{r}}\frac{\partial}{\partial k_{r}},\quad\mathsf{b}(z)=\sum_{r=1}^{n-1}\frac{1}{z-z_{r}}\bigg(\frac{\delta_{r}}{k_{r}}-k_{r}\frac{\partial^{2}}{\partial k_{r}^{2}}\bigg).$ (3.28)
We may observe that equations (3.18) imply
$\frac{\partial}{\partial y_{k}}F(\bm{k}(\bm{y},\rho))=\sum_{r=1}^{n-1}\frac{\partial k_{r}}{\partial y_{k}}\frac{\partial}{\partial k_{r}}F(\bm{k}(\bm{y},\rho))=\sum_{r=1}^{n-1}\frac{k_{r}}{y_{k}-z_{r}}\frac{\partial}{\partial k_{r}}F(\bm{k}(\bm{y},\rho)),$
yielding the key identities
${{\mathsf{a}}(y_{k})}=-\frac{\partial}{\partial y_{k}},\qquad k=1,\dots,n-3.$ (3.29)
Equations (3.29) furthermore imply
${{\mathsf{a}}^{2}(y_{k})}=\frac{\partial}{\partial y_{k}}{\mathsf{a}}(y_{k})=\frac{\partial^{2}}{\partial y_{k}^{2}}+{\mathsf{a}}^{\prime}(y_{k}).$ (3.30)
With the help of (3.27) and (3.30) it is now easy to see that
$\widetilde{\mathsf{T}}(y_{k})=-\frac{\partial^{2}}{\partial y_{k}^{2}},\qquad k=1,\dots,n-3.$ (3.31)
This is equivalent to (3.25).∎
3.3 The inverse transformation
Unitarity of the $\mathsf{SoV}$-transformation ensures existence of an inverse transformation. It can be realised explicitly by inverting the relation (3.19), taking advantage of the fact that equations (3.18) can be used to define functions $y_{k}=y_{k}(\bm{k})$ up to permutations $y_{k}\mapsto y_{l}$, $y_{l}\mapsto y_{k}$. The resulting map $\mathsf{SoV}^{-1}$ takes elements $\phi$ of ${\mathcal{H}}_{\rm Skl}$ to scaling-invariant functions of n2n-2 variables $\bm{k}=(k_{2},\dots,k_{n-1})$.
It next follows from (3.25) that $\mathsf{S}_{\bm{j},i}^{-1}$ maps elements $\phi_{\chi}$ of ${\mathcal{H}}_{\rm Skl}$ having the form
$\phi_{\chi}(\bm{y})=\prod_{k=1}^{n-3}\chi(y_{k},\bar{y}_{k}),\qquad\begin{aligned} &(\partial_{y}^{2}+t(y))\chi(y,\bar{y})=0,\\ &(\partial_{\bar{y}}^{2}+\bar{t}(\bar{y}))\chi(y,\bar{y})=0,\end{aligned}$ (3.32)
to eigenfunctions $\Psi_{\chi,i}$ of the differential operators $\mathsf{H}_{r}$, with eigenvalues $E_{r}$ determined by $t(y)$ using (2.9). One may furthermore easily check that $\phi_{\chi}(\bm{y})$ is normalizable, representing an element of ${\mathcal{H}}_{\rm Skl}$. This can be verified using (3.11), and taking into account that $\chi$ is analytic on $C$, having singular behaviour of the form
$\chi(y,\bar{y})=\sum_{\epsilon=\pm 1}|y-z_{r}|^{1+\epsilon(2j_{r}+1)}C_{r,\epsilon}(\bm{j})\big(1+{\mathcal{O}}(|y-z_{r}|\big),$ (3.33)
near the punctures $z_{r}$, $r=1,\dots,n$. Unitarity of the $\mathsf{SoV}$-transformation allows us to conclude that the eigenfunctions $\Psi_{\chi,i}=\mathsf{S}_{\bm{j},i}^{-1}(\phi_{\chi})$ are normalisable.
3.3.1 Smoothness
One should note, however, that the $L^{2}$-functions $\Psi_{\chi,i}=\mathsf{S}_{\bm{j},i}^{-1}(\phi_{\chi})$ are not necessarily point-wise defined, being defined by the Fourier-transformation of homogenous distributions.
Lemma 3.
The distributions $\Psi_{\chi,i}$ solve the holonomic system of differential equations
$\mathsf{H}_{r}\Psi_{\chi,i}=E_{r}\Psi_{\chi,i},\quad r=2,\dots,n-2,\qquad\sum_{r=2}^{n-1}(x_{r}\partial_{r}-j_{r}x_{r})\Psi_{\chi,i}=0,$
together with the complex conjugate equations.
Elliptic regularity implies that $\Psi_{\chi,i}$ can be represented by functions $\Psi_{\chi,i}(\bm{x})$ which are smooth away from the singularities of the differential operators $\mathsf{H}_{r}$. Such singularities are associated to the so-called wobbly bundles [EFK3].
However, one should note that the type of singular behaviour that can occur at the wobbly locus is strongly constrained by the unitarity of the $\mathsf{SoV}$ transformation. As $\Psi_{\chi,i}$ are in the image of ${\mathcal{H}}_{\rm Skl}$ under the unitary operators $\mathsf{S}_{\bm{j},i}^{-1}$, it is ensured that the singularities of $\Psi_{\chi,i}$ can not spoil square-integrability. The unitarity of the $\mathsf{SoV}$-transformation allows us to avoid a detailed analysis of the singularities that may occur at the wobbly loci.
3.3.2 Meromorphic continuation
Let us finally note another property of the inverse transformation $\mathsf{S}_{\bm{j},i}^{-1}(\phi_{\chi})$ that will become important below. Considering the dependence of the wave-functions $\Psi_{\chi,i}$ with respect to the parameters $j_{1},\dots,j_{n}$, one may note that this dependence is meromorphic. In order to see this, one may first note that the dependence of $\chi$ on $j_{1},\dots,j_{n}$ is analytic, as $\chi$ is the solution to a system of differential equations having coefficients depending analytically on $j_{1},\dots,j_{n}$. The regular distribution associated to the function $\widetilde{\Psi}_{\bm{E}}(\bm{k})$ defined by inverting the relation (3.19) also depends analytically on $j_{1},\dots,j_{n}$. The inverse Fourier-transformation $\mathsf{F}_{\bm{j},i}^{-1}$ is meromorphic in $J$, as can be seen by an argument very similar to the one in [GrS, Section 2].
We may therefore consider the meromorphic continuation of wave-functions $\Psi_{\chi,i}$ away from the range $j_{r}\in-\frac{1}{2}+\mathrm{i}{\mathbb{R}}$ in which the unitarity of the $\mathsf{SoV}$-transformation holds.
4 Outline of the proof of the analytic Langlands correspondence
Theorem 1 will follow from the following two results.
  • i)
    (Separation of variables) Any eigenstate $\Psi_{\chi}\in{\mathcal{H}}_{i}$ of ${\mathsf{H}}_{w}$ defines an element $\Phi_{\chi}=\mathsf{S}_{\bm{j},i}\Psi_{\chi}$ of ${\mathcal{H}}_{\rm\mathsf{SoV}}$. The wave-function $\Phi_{\chi}(\bm{y})$ representing $\Phi_{\chi}$ has the form
    $\Phi_{\chi}(\bm{y})=\prod_{k=1}^{n-3}\chi(y_{k},\bar{y}_{k}),$ (4.34)
    where $\chi(y,\bar{y})$ satisfies
    $(\partial_{y}^{2}+t_{\chi}(y))\chi(y,\bar{y})=0,\qquad(\bar{\partial}_{\bar{y}}^{2}+\bar{t}_{\chi}(\bar{y}))\chi({y},\bar{y})=0.$
    Single-valuedness of $\chi$ implies that $\partial_{y}^{2}+t_{\chi}(y)$ must be a real oper.
  • ii)
    (Hecke eigenvalue property) To each real oper $\partial_{w}^{2}+t(w)$ we may associate the single-valued solution $\chi$, unique up to scaling, of the pair of differential equations
    $(\partial_{w}^{2}+t(w))\chi(y,\bar{y})=0,\qquad(\bar{\partial}_{\bar{w}}^{2}+\bar{t}(\bar{w}))\chi(y,\bar{y})=0.$ (4.35)
    The functions $\mathsf{S}_{\bm{j},i}^{-1}\Phi_{\chi}$ associated to $\chi$, with $\Phi_{\chi}$ defined in (4.34), are eigenstates of ${\mathsf{Q}}_{w}$ with eigenvalue $\chi(w,\bar{w})$.
In order to prove these two results, we shall consider the $\mathsf{SoV}$ transformations $\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}$, associated to $C_{0,n+1}=\mathbb{P}^{1}\setminus\{z_{0},\dots,z_{n}\}$, $\bm{j^{\prime}}=(j_{0},j_{1},\dots,j_{n})$. The definition of $\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}$ is completely analogous to the definition of $\mathsf{S}_{\bm{j},i}^{(n)}:=\mathsf{S}_{\bm{j},i}$ given in the previous section, assigning index 0 to all variables associated to the additional puncture $z_{0}$, and admitting the factorisation $\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}=\mathsf{SoV}^{(n+1)}\circ\mathsf{F}_{\bm{j},i}^{(n+1)}$. In the special case $j_{0}=-1$, we will exhibit relations between $\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}$ and composition of $\mathsf{S}_{\bm{j},i}^{(n)}$ with the Hecke operators, implying the results i), ii) formulated above.
4.1 SoV-transform of eigenstates
The interplay between $\mathsf{SoV}$-transform and Hecke operators will prove one direction of the analytic Langlands correspondence rather easily.
Proposition 2.
The $\mathsf{SoV}$-transform $\Phi_{\chi}=\mathsf{S}_{\bm{j},i}^{(n)}\Psi_{\chi}$ of an eigenstate $\Psi_{\chi}$ can be represented in terms of the eigenvalue $\chi(z_{0},\bar{z}_{0})$ of the Baxter operators $\mathsf{H}_{z_{0}}$ as
$\Phi_{\chi}(\bm{y})=\prod_{k=1}^{n-3}\chi(y_{k},\bar{y}_{k}).$
Proof.
Let us note that
$(\mathsf{F}^{(n)}_{\bm{j},i}{\mathsf{H}}_{z_{0}}\Psi)(\bm{k})=\lim_{k_{0}\to 0}(\mathsf{F}^{(n+1)}_{\bm{j}^{\prime},i}\Psi^{\prime})(\bm{k}^{\prime}),\qquad\Psi^{\prime}(\bm{x}^{\prime}):=\mathbb{H}_{z_{0},x_{0}}\Psi(\bm{x}),$ (4.36)
using the notations $\bm{x}^{\prime}=(x_{0},\dots,x_{n-1})$ and $\bm{k}^{\prime}=(k_{0},\dots,k_{n-1})$. When $\Psi\equiv\Psi_{\chi}$ is an eigenstate of ${\mathsf{H}}_{z_{0}}$ with eigenvalue $\chi(z_{0},\bar{z}_{0})$, we may conclude from (4.36) that
$\chi(z_{0},\bar{z}_{0})(\mathsf{F}^{(n)}_{\bm{j},i}\Psi_{\chi})(\bm{k})=\lim_{k_{0}\to 0}(\mathsf{FT}^{(n+1)}_{\bm{j}^{\prime},i}\Psi^{\prime}_{\chi})(\bm{k}^{\prime}).$ (4.37)
The change of variables $\mathsf{SoV}^{(n+1)}$, explicitly given by
$k_{0}^{\prime}(\bm{y}^{\prime})=y\,(z_{0}-y_{0})\frac{\prod_{k=1}^{n-3}(z_{0}-y_{k})}{\prod_{r=1}^{n-1}(z_{0}-z_{s})},\quad k_{r}^{\prime}(\bm{y}^{\prime})=y\,\frac{z_{r}-y_{0}}{z_{r}-z_{0}}k_{r}(\bm{y}),\quad r=1,\dots,n.$ (4.38)
can be used to define the $\mathsf{SoV}$ transform $\phi^{\prime}=\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}\Psi^{\prime}_{\chi}$ of the function $\Psi^{\prime}_{\chi}$ as before. The function $\phi^{\prime}(y_{0},\dots,y_{n-3})$ is manifestly symmetric in $y_{0},\dots,y_{n-3}$.
Noting that $y_{0}\to z_{0}$ implies that $k_{0}^{\prime}(\bm{y}^{\prime})\to 0$ and $k_{r}^{\prime}(\bm{y}^{\prime})\to k_{r}(\bm{y})$, it follows from (4.37) that
$\lim_{y_{0}\to z_{0}}\big(\mathsf{S}^{(n+1)}_{\bm{j}^{\prime},i}\Psi^{\prime}_{\chi}\big)(\bm{y}^{\prime})=\chi(z_{0},\bar{z}_{0})\big(\mathsf{S}_{\bm{j},i}^{(n)}\Psi_{\chi}\big)(\bm{y}).$ (4.39)
As the function on the left side of (4.39) is symmetric under permutations of the arguments exchanging $z_{0}$ with $y_{k}$ for any $k=1,\dots,n-3$, it follows that
$\lim_{y_{0}\to z_{0}}\big(\mathsf{S}^{(n+1)}_{\bm{j}^{\prime},i}\Psi^{\prime}_{\chi}\big)(\bm{y}^{\prime})=\chi(z_{0},\bar{z}_{0})\prod_{k=1}^{n-3}\chi(y_{k},\bar{y}_{k}).$
The claim follows immediately. ∎
As the $\mathsf{SoV}$-transform can be inverted, we see that eigenstates are uniquely determined by their eigenvalues, which means that the spectrum is simple.
4.2 Hecke eigenvalue property
We have seen that eigenstates define real opers as their Hecke eigenvalues. It remains to prove that all real opers define eigenstates. To this aim, we shall consider inverse $\mathsf{SoV}$ transformations of the form $\big(\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}\big)^{-1}\phi^{\prime}_{\chi}$, where
$\phi^{\prime}_{\chi}(\bm{y}^{\prime})=\chi(y_{0},\bar{y}_{0})\phi_{\chi}(\bm{y}),\quad\phi_{\chi}(\bm{y})=\prod_{k=1}^{n-3}\chi(y_{k},\bar{y}_{k}),\quad\begin{aligned} &\bm{y}=(y_{1},\dots,y_{n-3}),\\ &\bm{y}^{\prime}=(y_{0},y_{1}\dots,y_{n-3}),\end{aligned}$ (4.40)
where $\chi(y,\bar{y})$ is the solution to the differential equation associated to a real oper on $C_{0,n}$ rather than $C_{0,n+1}$. Opers on $C_{0,n+1}$ are differential operators of the form $\partial_{y}^{2}+t^{(n+1)}(y)$, with
$t^{(n+1)}(y)=\sum_{r=0}^{n}\bigg(\frac{j_{r}(j_{r}+1)}{(y-z_{r})^{2}}+\frac{E_{r}}{y-z_{r}}\bigg).$ (4.41)
We see that opers on $C_{0,n+1}$ represent opers on $C_{0,n}$ if $j_{0}=-1$ and $E_{0}=0$.
The following theorem expresses a crucial link between the inverse Separation of Variables transformation and the Hecke modifications.
Theorem 3.
The Hecke modification of the states $\big(\mathsf{S}_{\bm{j},1}^{(n)}\big)^{-1}\phi_{\chi}\in{\mathcal{H}}_{1}$ can be represented as
$\big(\mathsf{S}_{\bm{j}^{\prime},0}^{(n+1)}\big)^{-1}\phi^{\prime}_{\chi}=\chi_{n}\,\frac{\pi\,\Gamma(-2j_{n})}{\Gamma(2j_{n}+1)}\,\mathbb{H}_{z_{0},x_{0}}\big(\mathsf{S}_{\bm{j},1}^{(n)}\big)^{-1}\phi_{\chi},$ (4.42)
using the notation $\chi_{n}:=C_{n,-}(\bm{j})$, referring to (3.33) for the definition of $C_{n,-}(\bm{j})$.
This theorem is proven in Section 6. It will be the key to one of our main results:
Theorem 4.
The states $\psi_{\chi}=\big(\mathsf{S}_{\bm{j},i}^{(n)}\big)^{-1}\phi_{\chi}\in{\mathcal{H}}_{i}$ are eigenstates of the operators $\mathsf{Q}_{y}$,
$\mathsf{Q}_{w}\,\psi_{\chi}=\frac{1}{\chi_{n}}\chi(w,\bar{w})\,\psi_{\chi}.$ (4.43)
The proof will also be given in Section 6. We shall next turn to a few preparations.
5 An alternative integral representation
As a preparation for the proof of Theorem 3, we will here consider alternative representations for the eigenfunctions which do not use the simplifying choice $z_{n}=\infty$.
5.1 Definitions
A main ingredient of the integrands will be a single-valued function $\chi(y,\bar{y})$ satisfying
$\begin{aligned} &(\partial_{y}^{2}+t(y))\chi(y,\bar{y})=0,\\ &(\partial_{\bar{y}}^{2}+\bar{t}(\bar{y}))\chi(y,\bar{y})=0,\end{aligned}\qquad t(y)=\sum_{r=1}^{n}\bigg(\frac{\delta_{r}}{(y-z_{r})^{2}}+\frac{E_{r}}{y-z_{r}}\bigg).$ (5.44)
in both cases. We shall consider the following integrals
$\Theta^{(n,i)}_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[3.22916pt] x_{1}&\dots&x_{n}\end{smallmatrix}\Big)=\frac{1}{\pi^{n}}\int\prod_{r=1}^{n}\frac{d^{2}k_{r}\,e^{2\mathrm{i}\,\mathrm{Im}(k_{r}x_{r})}}{|k_{r}|^{2j_{r}+2}}\prod_{a=0}^{i}\delta^{(2)}\big({\textstyle\sum_{r=1}^{n}z_{r}^{a}k_{r}}\big)$ (5.45)
$\hskip 184.9429pt\times|\hat{\rho}_{i}(\bm{k})|^{2}\prod_{k=1}^{n-2-i}\chi\big(y_{k}(\bm{k}),\bar{y}_{k}(\bm{k})\big),$
with $y_{1}(\bm{k}),\dots,y_{n-2-i}(\bm{k})$ and $\hat{\rho}_{i}(\bm{k})$, $\bm{k}=(k_{1},\dots,k_{n})$, defined by
$k_{r}=\hat{\rho}_{i}(\bm{k})\frac{\prod_{k=1}^{n-2-i}(z_{r}-y_{k}(\bm{k}))}{\prod_{s\neq r}(z_{r}-z_{s})},\qquad r=1,\dots,n,\qquad\hat{\rho}_{i}(\bm{k})=\sum_{r=1}^{n}z_{r}^{1+i}k_{r},$ (5.46)
for $i=0,1$, respectively. The main differences between $i=0$ and $i=1$ are a larger range of values for the indices $y_{k}$ in the case $i=0$, and the extra delta-distribution $\delta^{(2)}\big({\textstyle\sum_{r=1}^{n}z_{r}k_{r}}\big)$, in the case $i=1$.
The functions $\Theta^{(n,i)}_{\chi}$ will turn out to be proportional to the functions $\Sigma^{(n,i)}_{\chi}\big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[1.07639pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\big)$ defined by $\big(\mathsf{S}_{\bm{j},i}^{(n)}\big)^{-1}\phi_{\chi}$ in the limit $z_{n}\to\infty$.
The integrals on the right side of (LABEL:Fourier(n)) are initially defined in the distributional sense. In order to see that (LABEL:Fourier(n)) actually defines smooth functions away from singular divisors of co-dimension one, we may consider the inverse $\mathsf{SoV}$-transformation in the case where the number $n$ of punctures is replaced by $m=n+1$, with punctures $z_{r}$ now labelled by an index $r$ running from $1$ to $n+1$, assuming $z_{n+1}=\infty$. The inverse $\mathsf{SoV}$-transformation will in this case act on products of functions $\hat{\chi}$ representing real opers on the $(n+1)$-punctured sphere. As noted above, it will define a function denoted as $\Sigma^{(n+1)}_{\hat{\chi}}\big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[2.15277pt] x_{1}&\dots&x_{n}\end{smallmatrix};j_{n+1}\big)$ away from singular divisors of co-dimension one. We furthermore know that $\Sigma^{(n+1)}_{\hat{\chi}}$ is meromorphic in $j_{n+1}$, with poles depending only on $j_{n+1}-\sum_{r=1}^{n}j_{r}$. For generic values of $j_{1},\dots,j_{n}$ we may therefore consider the limit $j_{n+1}\to 0$, combined with $E_{n+1}\to 0$, yielding the right side of (LABEL:Fourier(n)).
5.2 Differential equations
We are now going to see that $\Theta^{(n,i)}_{\chi}$ satisfy a holonomic system of differential equations identifying $\Theta^{(n,i)}_{\chi}$ as representations of the eigenfunctions with $z_{n}$ in generic position, and an additional variable $x_{n}$, for $i=0,1$. We will see that the dependence $\Theta^{(n,i)}_{\chi}\big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[2.15277pt] x_{1}&\dots&x_{n}\end{smallmatrix}\big)$ with respect to the variables $x_{1},\dots,x_{n}$ is constrained by certain groups of symmetries extending the symmetry of $\Sigma^{(n,i)}_{\chi}\big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[2.15277pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\big)$ under translations and dilatations. Invariance under the Lie-algebras of the symmetry groups is contained in the following Lemma.
Lemma 4.
The functions $\Theta^{(n,i)}_{\chi}$ defined in (LABEL:Fourier(n)) satisfy the system of equations
$\mathsf{H}_{r}\Theta^{(n,i)}_{\chi}=E_{r}\Theta^{(n,i)}_{\chi},\qquad{\mathcal{D}}^{a}\Theta^{(n,i)}_{\chi}=0,\quad\begin{aligned} &a\in\{0,+\},\\ &i\in\{0,1\},\end{aligned}\qquad\left\{\begin{aligned} &{\mathcal{D}}^{-}\Theta^{(n,0)}_{\chi}=0,\\ &\widehat{\mathcal{D}}^{+}\Theta^{(n,1)}_{\chi}=0,\end{aligned}\right$. (5.47)
where $r\in\{1,\dots,n\}$, $a\in\{-,0,+\}$ and $i\in\{0,1\}$, using the notations
$\mathcal{D}^{0}=\sum_{r=1}^{n}\bigg(x_{r}\frac{\partial}{\partial x_{r}}-j_{r}\bigg)\,,\quad\mathcal{D}^{-}=\sum_{r=1}^{n}\bigg(x_{r}^{2}\frac{\partial}{\partial x_{r}}-2j_{r}x_{r}\bigg),\quad{\mathcal{D}}^{+}=\sum_{r=1}^{n}\frac{\partial}{\partial x_{r}}\,$, (5.48)
$\mathsf{H}_{r}=\sum_{s\neq r}^{n}\frac{{\mathcal{D}}_{r}^{a}{\mathcal{D}}_{s}^{b}\eta_{ab}}{z_{r}-z_{s}},\qquad r=1,\dots,n,\hskip 105.2751pt\widehat{\mathcal{D}}^{+}=\sum_{r=1}^{n}z_{r}\frac{\partial}{\partial x_{r}}.$ (5.49)
Proof.
For the following argument we shall again start by considering $\Sigma^{(n+1)}_{\hat{\chi}}\big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[2.15277pt] x_{1}&\dots&x_{n}\end{smallmatrix};j_{n+1}\big)$, as done at the beginning of this section. $\Sigma^{(n+1)}_{\hat{\chi}}$ satisfies the system of differential equations,
$\mathsf{H}_{r}\Sigma^{(n+1)}_{\chi}=E_{r}\Sigma^{(n+1)}_{\chi},\quad r=1,\dots,{n},\qquad{\mathcal{D}}_{a}\Sigma^{(n+1)}_{\chi}=0,\qquad a\in\{0,+\},$ (5.50)
as follows from Proposition 1. We are going to note that the differential equations stated in Lemma 4 follow from (5.50) in the limit $j_{n+1}\to 0$ and $E_{m+1}\to 0$. The only nontrivial case is the equation ${\mathcal{D}}_{-}\Theta^{(n,0)}_{\chi}=0$.
Let us consider the generating function of differential operators
$\mathsf{T}_{n+1}(z)=\sum_{r=1}^{n}\bigg(\frac{\delta_{r}}{(y-z_{r})^{2}}+\frac{{\mathsf{H}}_{r}}{y-z_{r}}\bigg).$
The Fourier-transformation $\widetilde{\mathsf{T}}_{n+1}(z)$ of $\mathsf{T}_{n+1}(z)$ can be represented in the form (3.27) with $\mathsf{a}(z)$, $\mathsf{b}(z)$ and $\mathsf{c}(z)$ replaced by very similar differential operators, defined by extending the summation over $r$ in the formulae (3.28) to include the value $r=n$. We may note that
$\widetilde{\mathsf{T}}_{n+1}(z)$ $=\frac{1}{z^{2}}(\mathsf{a}_{\infty}^{2}-\mathsf{a}_{\infty}+\mathsf{c}_{\infty}\mathsf{b}_{\infty})+\frac{\rho(\bm{k})}{z^{3}}\,\mathsf{b}_{\infty}+{\mathcal{O}}(z^{-4})$ (5.51)
$=\frac{\delta_{n+1}}{z^{2}}+\frac{E_{n+1}}{z^{3}}+{\mathcal{O}}(z^{-4}),$
using the notations
$\mathsf{a}_{\infty}=-\sum_{r=1}^{n}k_{r}\frac{\partial}{\partial k_{r}},\qquad\mathsf{b}_{\infty}=\sum_{r=1}^{n}\bigg(\frac{\delta_{r}}{k_{r}}-k_{r}\frac{\partial^{2}}{\partial k_{r}^{2}}\bigg),\qquad\mathsf{c}_{\infty}=\sum_{r=1}^{n}k_{r}.$
As the Fourier-transformation defining ${\Theta}^{(n)}_{\chi}$ is supported on $\sum_{r=1}^{n}k_{r}=0$, and as we have $\mathsf{a}_{\infty}\widetilde{\Theta}^{(n)}_{\chi}=j_{n+1}$ by dilatation invariance, we may conclude from (5.51) that $\widetilde{\Theta}^{(n)}_{\chi}=0$ must satisfy $\mathsf{b}_{\infty}\widetilde{\Theta}^{(n)}_{\chi}=0$ in the limit $j_{n+1}\to 0$ and $E_{n+1}\to 0$. This implies the equation ${\mathcal{D}}_{-}\Theta^{(n)}_{\chi}=0$ by Fourier-transformation. ∎
5.3 Relation to inverse SoV-transformation
We will now show that the functions $\Sigma^{(n,i)}_{{\chi}}$ representing the inverse $\mathsf{SoV}$-transformations can be represented as a certain limit of the functions $\Theta^{(n,i)}_{\chi}$ defined in (LABEL:Fourier(n)), for $i=0,1$, respectively. The integrals defining $\Theta^{(n,i)}_{\chi}$ thereby furnish an alternative representation for the inverse $\mathsf{SoV}$-transformations that will be used in the next section.
Lemma 5.
$\lim_{z_{n}\to\infty}|{z_{n}}|^{-2j_{n}}\Theta^{(n,1)}_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[4.09024pt] x_{1}&\dots&x_{n}\end{smallmatrix}\Big)=\Sigma^{(n,1)}_{{\chi}}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[4.09024pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big).$ (5.52)
Proof.
We will use the distribution $\delta^{(2)}\big({\textstyle\sum_{r=1}^{n}z_{r}k_{r}}\big)$ in order to perform the integration over $k_{n}$, using the fact that $\sum_{r=1}^{n-1}k_{r}z_{r}$ stays finite in the limit of our interest. It follows that we have
$k_{n}=-z_{n}^{-1}\rho(\bm{k}),\quad\text{and}\quad\hat{\rho}_{1}(\bm{k})=-{z_{n}}\rho(\bm{k})+{\mathcal{O}}(z_{n}^{0}),$ (5.53)
on the support of $\delta^{(2)}\big({\textstyle\sum_{r=1}^{n}z_{r}k_{r}}\big)$, recalling that $\rho(\bm{k})=\sum_{r=1}^{n-1}k_{r}z_{r}$, and $\hat{\rho}_{1}(\bm{k})=\sum_{r=1}^{n}k_{n}z_{n}^{2}$. The relations defining the functions $y_{k}(\bm{k})$ simplify to
$k_{r}=\frac{\hat{\rho}_{1}(\bm{k})}{(z_{r}-z_{n})}\frac{\prod_{k=1}^{n-3}(z_{r}-y_{k}(\bm{k}))}{\prod_{r\neq s}^{n-1}(z_{r}-z_{s})}=\rho(\bm{k})\frac{\prod_{k=1}^{n-3}(z_{r}-y_{k}(\bm{k}))}{\prod_{r\neq s}^{n-1}(z_{r}-z_{s})}+\mathcal{O}(z^{-1}_{n}),$ (5.54)
for $r=1,\dots,n-1$. We can thereby conclude that the functions $y_{k}(\bm{k})$ become independent of $k_{n}$ in the limit $z_{n}\to\infty$. In this way it becomes straightforward to find that the limit of the integral defining $\Theta^{(n,1)}_{\chi}$ on the right side of (5.52) is given by the integral
$\frac{1}{\pi^{n}}\int\prod_{r=1}^{n-1}\frac{d^{2}k_{r}\,e^{2\mathrm{i}\,\mathrm{Im}(k_{r}x_{r})}}{|k_{r}|^{2j_{r}+2}}\delta^{(2)}\big({\textstyle\sum_{r=1}^{n}k_{r}}\big)|\rho(\bm{k})|^{-2j_{n}-2}\prod_{k=1}^{n-3}\chi\big(y_{k}(\bm{k}),\bar{y}_{k}(\bm{k})\big).$ (5.55)
This is the integral defining $\Sigma^{(n,1)}_{{\chi}}\big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[1.07639pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\big)$. ∎
For future use, let us note an interesting consequence of (5.55): We have
$\lim_{z_{n}\to\infty}\Theta^{(n,1)}_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}&-1\\[4.09024pt] x_{1}&\dots&x_{n-1}&x_{n}\end{smallmatrix}\Big)=\Theta^{(n-1,0)}_{{\chi}}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[4.09024pt] x_{1}&\dots&x_{n-1}\end{smallmatrix}\Big).$ (5.56)
This relation will play an important role in our proof of Theorem 3 later.
Proposition 3.
We have
$\lim_{x_{n},z_{n}\to\infty}\Big|\frac{z_{n}}{x_{n}}\Big|^{4j_{n}}\Theta^{(n,0)}_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[4.09024pt] x_{1}&\dots&x_{n}\end{smallmatrix}\Big)=\chi_{n}\,\frac{\pi\,\Gamma(-2j_{n})}{\Gamma(2j_{n}+1)}\Sigma^{(n,0)}_{{\chi}}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[4.09024pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big),$ (5.57)
with $\chi_{n}$ being a short notation for $C_{n,-}(\bm{j})$ defined in (3.33).
Proof.
In order to verify (5.57) let us note that the leading contributions to the limit $x_{n}\to\infty$ are coming from the region where $k_{n}\to 0$. It follows from (5.46) that there must exist an index $k$ such that $y_{k}\to z_{n}$. The permutation symmetry of the integrand allows us to assume without loss of generality that $y_{n-2}\to z_{n}$. We may next note that
$k_{r}=\rho\frac{(z_{r}-y_{n-2})}{(z_{r}-z_{n})}\frac{\prod_{k=1}^{n-3}(z_{r}-y_{j})}{\prod_{r\neq s}^{n-1}(z_{r}-z_{s})}=\rho\frac{\prod_{k=1}^{n-3}(z_{r}-y_{k})}{\prod_{r\neq s}^{n-1}(z_{r}-z_{s})}+\mathcal{O}(x^{-1}_{n}).$ (5.58)
It follows that the functions $y_{k}(\bm{k})$, $k=1,\dots,n-3$, and $\hat{\rho}_{0}(\bm{k})$ become independent of $k_{n}$ in the limit of our interest.
In the limit $y_{n-2}\to z_{n}$ we may furthermore use
$k_{n}=\rho{(z_{n}-y_{n-2})}\frac{\prod_{k=1}^{n-3}(z_{n}-y_{j})}{\prod_{s=1}^{n-1}(z_{n}-z_{s})},$ (5.59)
together with the asymptotic behaviour (3.33). As $k_{n}$ is proportional to $y_{n-2}-z_{n}$, we see that the dependence of the integrand defining the right side of (5.57) with respect to $k_{n}$ simplifies significantly. We may factorise the Fourier-transformation over $k_{1},\dots,k_{n}$ into an integration over $k_{n}$, followed by a Fourier-transformation over $k_{1},\dots,k_{n-1}$. The integration over $k_{n}$ can be carried out explicitly using the identity (see e.g. equation (7) of Appendix B1.7 in [GeS])
$\frac{2j+1}{\pi}\int_{\mathbb{C}}d^{2}x^{\prime}\;|x-x^{\prime}|^{-4j-4}e^{kx^{\prime}-\bar{k}\bar{x}^{\prime}}=e^{kx-\bar{k}\bar{x}}|k|^{2(2j+1)}\frac{\Gamma(-2j-1)}{\Gamma(2j+1)}\,.$ (5.60)
One of the resulting terms is a distribution supported at $x_{n}=0$ which can not contribute to the limit $x_{n}\to\infty$. The other term has an asymptotic behaviour for $z_{n}\to\infty$ of the following form
$\frac{\chi_{n}}{\pi^{n}}\left|\frac{x_{n}}{z_{n}}\right|^{4j_{n}}\frac{\pi\,\Gamma(-2j_{n})}{\Gamma(2j_{n}+1)}\int\prod_{r=1}^{n-1}\frac{d^{2}k_{r}\,e^{2\mathrm{i}\,\mathrm{Im}(k_{r}x_{r})}}{|k_{r}|^{2j_{r}+2}}\delta^{(2)}\big({\textstyle\sum_{r=1}^{n-1}}k_{r}\big)|\rho(\bm{k})|^{2(j_{n}+1)}\prod_{k=1}^{n-3}\chi\big(y_{k}(\bm{k}),\bar{y}_{k}(\bm{k})\big),$
which is proportional to the left side of (5.57). ∎
By solving the differential equations satisfied by $\Theta^{(n,i)}_{\chi}$ with the boundary condition provided by (5.57) one may express the functions $\Theta^{(n)}_{\chi}$ defined in (LABEL:Fourier(n)) in terms of $\Sigma^{(n)}_{\chi}\big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[2.15277pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\big)$.
6 Hecke modifications of the inverse SoV transformation
We shall here prove the crucial Theorems 3 and 4. We will consider the inverse $\mathsf{SoV}$-transformations associated to Riemann spheres $C_{0,n+1}$ obtained from $C_{0,n}$ by adding another marked point at $z=z_{0}$. When the parameter $j_{0}$ associated to $z_{0}$ is chosen as $j_{0}=-1$, we will find a relation with the Hecke modification $\mathbb{H}_{z_{0},x_{0}}$, and with the Hecke operator $\mathsf{H}_{z_{0}}$.
6.1 Differential equations
It will be crucial to observe that the system of differential equations satisfied by $\Theta^{(n+1,1)}_{\chi}$ simplifies somewhat when $j_{0}=-1$, and $E_{0}=0$.
Proposition 4.
When $j_{0}=-1$ and $E_{0}=0$, the distribution $\Theta^{(n+1,1)}_{\chi}$ satisfies the differential equations ${\mathcal{D}}_{a}\Theta^{(n+1,1)}_{\chi}=0$, $a\in\{-,0,+\}$, and furthermore
$\mathcal{N}\Theta^{(n+1,1)}_{\chi}=0,\qquad\mathcal{N}=\sum_{r=1}^{n}\frac{1}{z_{0}-z_{r}}\bigg((x_{r}-x_{0})^{2}\frac{\partial}{\partial x_{r}}-2j_{r}(x_{r}-x_{0})\bigg).$ (6.61)
Proof.
The condition $E_{0}=0$ implies that
[…]
with $\widetilde{\Theta}^{(n+1,1)}_{\chi}$ being the Fourier-transformation of ${\Theta}^{(n+1,1)}_{\chi}$, with $\widetilde{\mathsf{T}}(z)$ defined in (3.27), and
$\widetilde{\mathcal{N}}:=\sum_{r=1}^{n-1}\frac{k_{r}}{z_{0}-z_{r}}\bigg(\frac{\partial^{2}}{\partial k_{r}^{2}}-\frac{\delta_{r}}{k_{r}^{2}}-2\frac{\partial^{2}}{\partial k_{r}\partial k_{0}}+\frac{\partial^{2}}{\partial k_{0}^{2}}\bigg).$
It follows that $\widetilde{\mathcal{N}}\widetilde{\Theta}^{(n+1,1)}_{\chi}=0$, implying (6.61) by inverse Fourier-transformation. ∎
Remark 1.
Equation (6.61) can be recognised as the differential equation characterising Hecke modifications of the conformal blocks of affine Lie algebras [AFW].
6.2 Proof of Theorems 3 and 4
Proof.
(Theorem 3) We consider the integrals $\Theta^{(n+1,1)}_{\chi}$ defined in (LABEL:Fourier(n)) in the special cases where $j_{0}=-1$ and $E_{0}=0$. With the help of the system of differential equations stated in Proposition 4 we may show that there exists a function $\Psi_{\chi}\big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[1.07639pt] x_{1}&\dots&x_{n}\end{smallmatrix}\big)$ satisfying
$\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[1.07639pt] \frac{ax_{1}-b}{cx_{1}-d}&\dots&\frac{ax_{n}-b}{cx_{n}-d}\end{smallmatrix}\Big)=\prod_{r=1}^{n}\frac{1}{|cx_{r}+d|^{4j_{r}}}\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[3.22916pt] x_{1}&\dots&x_{n}\end{smallmatrix}\Big),$
which allows us to represent $\Theta^{(n+1,1)}_{\chi}$ in the form
$\Theta^{(n+1,1)}_{\chi}\Big(\begin{smallmatrix}j_{0}&\dots&j_{n}\\[3.22916pt] x_{0}&\dots&x_{n}\end{smallmatrix}\Big)=\prod_{r=1}^{n}\frac{|x_{r}-x_{0}|^{4j_{r}}}{|z_{r}-z_{0}|^{2j_{r}}}\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[1.07639pt] -\frac{z_{1}-z_{0}}{x_{1}-x_{0}}&\dots&-\frac{z_{n}-z_{0}}{x_{n}-x_{0}}\end{smallmatrix}\Big).$ (6.62)
We may next note that
$\lim_{z_{0}\to\infty}\prod_{r=1}^{n}\frac{|x_{r}-x_{0}|^{4j_{r}}}{|z_{r}-z_{0}|^{2j_{r}}}\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[1.07639pt] -\frac{z_{1}-z_{0}}{x_{1}-x_{0}}&\dots&-\frac{z_{n}-z_{0}}{x_{n}-x_{0}}\end{smallmatrix}\Big)=\lim_{z_{0}\to\infty}\prod_{r=1}^{n}\frac{|x_{r}-x_{0}|^{4j_{r}}}{|z_{0}|^{2j_{r}}}\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[1.07639pt] \frac{z_{0}}{x_{1}-x_{0}}&\dots&\frac{z_{0}}{x_{n}-x_{0}}\end{smallmatrix}\Big)$
$\hskip 156.49014pt=\prod_{r=1}^{n}{|x_{r}-x_{0}|^{4j_{r}}}\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[1.07639pt] \frac{1}{x_{1}-x_{0}}&\dots&\frac{1}{x_{n}-x_{0}}\end{smallmatrix}\Big)=\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[3.22916pt] x_{1}&\dots&x_{n}\end{smallmatrix}\Big).$
The same limit of the left side of (6.62) can be computed with the help of equation (5.56). We conclude that
$\Psi_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[3.22916pt] x_{1}&\dots&x_{n}\end{smallmatrix}\Big)=\Theta^{(n,0)}_{{\chi}}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[4.09024pt] x_{1}&\dots&x_{n}\end{smallmatrix}\Big),$ (6.63)
and therefore
$\Theta^{(n+1,1)}_{\chi}\Big(\begin{smallmatrix}-1&j_{1}&\dots&j_{n}\\[3.22916pt] x_{0}&x_{1}&\dots&x_{n}\end{smallmatrix}\Big)=\prod_{r=1}^{n}\frac{|x_{r}-x_{0}|^{4j_{r}}}{|z_{r}-z_{0}|^{2j_{r}}}\Theta^{(n,0)}_{{\chi}}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n}\\[1.07639pt] -\frac{z_{1}-z_{0}}{x_{1}-x_{0}}&\dots&-\frac{z_{n}-z_{0}}{x_{n}-x_{0}}\end{smallmatrix}\Big).$ (6.64)
Theorem 3 now follows from (6.64) by taking the limit $z_{n}\to\infty$, $x_{n}\to\infty$ using equations (5.52) and (5.57), leading to
$\Sigma^{(n+1,1)}_{\chi}\Big(\begin{smallmatrix}-1&j_{1}&\dots&j_{n-1}\\[3.22916pt] x_{0}&x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big)=\;\chi_{n}\;\frac{\pi\,\Gamma(-2j_{n})}{\Gamma(2j_{n}+1)}\;\Sigma^{(n,0)}_{{\chi}}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[1.07639pt] -\frac{z_{1}-z_{0}}{x_{1}-x_{0}}&\dots&-\frac{z_{n-1}-z_{0}}{x_{n-1}-x_{0}}\end{smallmatrix};j_{n}\Big),$ (6.65)
which is equivalent to Theorem 3. ∎
Proof.
(Theorem 4) In order to prove Theorem 4 using Theorem 3 it will help to note that the Hecke operator ${\mathsf{H}}_{z_{0}}$ can be represented in terms of the limit $j_{0}\to-1$, $x_{0}\to 0$ of the operator
$({\mathsf{R}}_{j_{0}}\Psi^{\prime})(\bm{x}^{\prime})=-\frac{2j_{0}+1}{\pi}\int_{\mathbb{C}}d^{2}x_{0}\;|x_{0}-x_{0}^{\prime}|^{-4j_{0}-4}{\Psi}^{\prime}(\bm{x}^{\prime}).$
Equation (6.65) implies
$\chi_{n}\;\frac{\pi\,\Gamma(-2j_{n})}{\Gamma(2j_{n}+1)}\;\mathsf{H}_{z_{0}}\Sigma^{(n,0)}_{{\chi}}=\lim_{j_{0}\to-1}{\mathsf{R}}_{j_{0}}\Sigma^{(n+1,1)}_{\chi}$. (6.66)
In order to compute the right side we will use the following proposition.
Proposition 5.
The inverse $\mathsf{SoV}$-transformation $\big(\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}\big)^{-1}$ satisfies the reflection property
${\mathsf{R}}_{j_{0}}\big(\mathsf{S}_{\bm{j}^{\prime},i}^{(n+1)}\big)^{-1}=\frac{\Gamma(-2j_{0})}{\Gamma(2j_{0}+2)}\big(\mathsf{S}_{\bm{j}^{\prime\prime},i}^{(n+1)}\big)^{-1},$ (6.67)
where $\bm{j}^{\prime\prime}=(-j_{0}-1,j_{1},\dots)$ if $\bm{j}^{\prime}=(j_{0},j_{1},\dots)$.
Proof.
(Proposition 5) The identity (5.60) implies ${\mathsf{R}}_{j_{0}}\big(\mathsf{F}_{\bm{j}^{\prime},i}^{(n+1)}\big)^{-1}=\frac{\Gamma(-2j)}{\Gamma(2j+2)}\big(\mathsf{F}_{\bm{j}^{\prime\prime},i}^{(n+1)}\big)^{-1}$, which immediately yields Proposition 5. ∎
We thereby find that
$\chi_{n}\;\frac{\pi\,\Gamma(-2j_{n})}{\Gamma(2j_{n}+1)}\;\big(\mathsf{H}_{z_{0}}\Sigma^{(n,0)}_{{\chi}}\big)\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[3.22916pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big)=$ (6.68)
$\qquad\qquad\qquad=\lim_{j_{0}\to-1}\frac{\Gamma(-2j_{0})}{\Gamma(2j_{0}+2)}\Sigma^{(n+1,1)}_{\chi}\Big(\begin{smallmatrix}-j_{0}-1&j_{1}&\dots&j_{n-1}\\[3.22916pt] x_{0}&x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big).$
Let us next observe that the identity
$\lim_{j\to-1}\frac{\Gamma(-2j-1)}{\Gamma(2j+2)}|y|^{2j}=\pi\,\delta^{2}(y),$
implies that
$\lim_{j_{0}\to-1}\frac{\Gamma(-2j_{0})}{\Gamma(2j_{0}+2)}\Sigma^{(n+1,1)}_{\chi}\Big(\begin{smallmatrix}-j_{0}-1&j_{1}&\dots&j_{n-1}\\[3.22916pt] x_{0}&x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big)=\pi\,\chi(z_{0},\bar{z}_{0})\,\Sigma^{(n,1)}_{\chi}\Big(\begin{smallmatrix}j_{1}&\dots&j_{n-1}\\[3.22916pt] x_{1}&\dots&x_{n-1}\end{smallmatrix};j_{n}\Big).$
Inserting this into (6.68) yields
$\mathsf{H}_{z_{0}}\Sigma^{(n,0)}_{{\chi}}=\frac{\Gamma(2j_{n}+1)}{\Gamma(-2j_{n})}\,\frac{\chi(z_{0},\bar{z}_{0})}{\chi_{n}}\,\Sigma^{(n,1)}_{\chi}$. (6.69)
Note, on the other hand, that
$\mathsf{H}_{n}\Psi_{1}:=\lim_{z_{0}\to z_{n}}|z_{r}-z_{0}|^{2j_{r}}\mathsf{H}_{z_{0}}\Psi_{1}=\frac{\Gamma(2j_{n}+1)}{\Gamma(-2j_{n})}\mathsf{H}_{\infty}\Psi_{1},$
for all $\Psi_{1}\in{\mathcal{H}}_{1}$. Taken together, we therefore find that
$\mathsf{Q}_{z_{0}}\,\Sigma^{(n,1)}_{{\chi}}=\mathsf{H}_{z_{0}}(\mathsf{H}_{n})^{-1}\,\Sigma^{(n,1)}_{{\chi}}=\frac{1}{\chi_{n}}\chi(z_{0},\bar{z}_{0})\;\Sigma^{(n,1)}_{\chi}$.
which is the Hecke eigenvalue property satisfied by $\Sigma^{(n,1)}_{\chi}$. ∎
Acknowledgements. The question about the unitarity of the $\mathsf{SoV}$-transformation was raised by D. Kazhdan in 2023 in a discussion with P. Etingof, B. Feigin, E. Frenkel, A. Goncharov, and one of the authors (J.T.). J.T. would like to thank all participants for these discussions. We’d in particular like to thank P. Etingof for pointing out reference [GrS] to us.
We’d furthermore like to acknowledge the support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy - EXC 2121 “Quantum Universe” - 390833306 and the Collaborative Research Center - SFB 1624 “Higher structures, moduli spaces and integrability” - 506632645.